### 30 Nov 2023 v1

# Accurate Segmentation of Optic Disc and Cup from Multiple Pseudo-labels by Noise-Aware Learning

Abstract
Optic disc and cup segmentation play a crucial role in automating the screening and diagnosis of optic glaucoma. While data-driven convolutional neural networks (CNNs) show promise in this area, the inherent ambiguity of segmenting object and background boundaries in the task of optic disc and cup segmentation leads to noisy annotations that impact model performance. To address this, we propose an innovative label-denoising method of Multiple Pseudo-labels Noise-aware Network (MPNN) for accurate optic disc and cup segmentation. Specifically, the Multiple Pseudo-labels Generation and Guided Denoising (MPGGD) module generates pseudo-labels by multiple different initialization networks trained on true labels, and the pixel-level consensus information extracted from these pseudo-labels guides to differentiate clean pixels from noisy pixels. The training framework of the MPNN is constructed by a teacher-student architecture to learn segmentation from clean pixels and noisy pixels. Particularly, such a framework adeptly leverages (i) reliable and fundamental insights from clean pixels and (ii) the supplementary knowledge within noisy pixels via multiple perturbation-based unsupervised consistency. Compared to other label-denoising methods, comprehensive experimental results on the RIGA dataset demonstrate our method’s excellent performance and significant denoising ability. The code is available at https://github.com/wwwtttjjj/MPNN
Index Terms:
Optic Disc and Cup Segmentation, Label-denoising, Multiple Pseudo-labels
I Introduction
Optic disc and cup segmentation is an important step in the automated screening and diagnosis of optic nerve head abnormalities such as glaucoma. With the rapid development of deep learning, various models based on convolutional neural networks (CNNs) have emerged, showcasing promising performance in medical image segmentation. These methods [1, 2, 3, 4, 5] leverage their inherent capacity to learn intricate patterns and features from extensive labeled data, enabling them to achieve high levels of accuracy in segmenting medical images. However, due to the inherent ambiguity of the true segmentation boundaries in optic disc and cup segmentation [6, 7] and the inevitable errors in manual labeling, the annotation often contains noise. Numerous studies [8, 9, 10, 11] have indicated that label noise can significantly impact the accuracy of models.
In order to mitigate the impact of label noise on medical image segmentation, a strategy is to use multiple expert annotations. Some progress has been achieved in the joint annotation learning from multiple experts for the task of optic disc and optic cup segmentation. Ji et al. [12] introduced MRNet, an innovative technique that capitalizes on the specialized insights of individual raters, treating their expertise as prior knowledge to extract high-level semantic features. Their proposed approach involves reconstructing multi-rater rankings based on initial predictions, subsequently harnessing the varying degrees of agreement among multiple raters to enhance the overall segmentation performance. By amalgamating expert knowledge and effectively pooling input from multiple raters, MRNet presents a promising avenue for improving the precision of optic disc and cup segmentation. Self-Calib [13] attempts to find ground truths by running divergent and convergent models in loops. The latter category of methods demonstrates a greater capacity for exploiting information from multiple annotations, resulting in a more impartial segmentation model.
However, it is worth noting that obtaining segmentation annotations from experts is a challenging task, not to mention the need for annotations from multiple experts. Therefore, it is necessary to explore other available methods.
Without introducing additional expert labels, a series of label-denoising strategies have been proposed to cope with the challenges posed by noisy labels in segmentation tasks. These strategies typically involve two steps: first, noisy labels are identified using various methods, and then pseudo-labels are used to refine these noisy labels or to reject them. Some methods [8, 14, 15] use Confident Learning [16] modules to find noisy labels and use refinement module for processing. The method based on Confidence Learning assesses the classification confidence of each pixel’s output from the network and identifies potential noisy pixels based on this confidence level. Subsequently, the refinement module uses pseudo-labels to replace these noisy labels. Other methods [17, 18, 19, 20, 21, 22, 23] involve training multiple networks and selecting labels with the minimum loss consensus between network predictions for training. In essence, these selected pixels are treated as clean pixels, thereby eliminating noisy pixels. Compared to multi-annotation learning, the direct denoising approach is simple, efficient, and does not require the introduction of additional expert knowledge. However, while these methods have shown promise in various domains, their adaptation and implementation for the specific task of optic disc and cup segmentation still require careful investigation.
In this paper, we propose an innovative label-denoising method of Multiple Pseudo-labels Noise-aware Network (MPNN) for accurate optic disc and cup segmentation. To address the issue of label noise, we propose the Multiple Pseudo-labels Generation and Guided Denoising (MPGGD) module. We use multiple networks with different initialization to fit the same training set, and after reaching a certain threshold stop training to obtain multiple pseudo-labels for the training set. Subsequently, the pixels corresponding to the labels that perform consistently on all the pseudo-labels are considered as clean pixels and the rest are considered as noisy pixels. Furthermore, the proposed MPNN consists of a teacher-student architecture, where the student network learns by (i) minimizing the segmentation loss on the set of clean pixels and (ii) minimizing the consistency loss on the set of noisy pixels in the teacher model with multiple uncertainties. The teacher network is updated based on the parameters of the student network using the Exponential Moving Average (EMA). Our contributions are as follows:
  • $\bullet$
    To the best of our knowledge, the proposed MPNN is the first method to improve the accuracy of the optic disc and cup from the perspective of label denoising.
  • $\bullet$
    We proposed the MPGGD as the first execution module. It separates reliable and unreliable information, provides accurate priors for the segmentation network, and avoids the negative impact of label noise on the network.
Figure 1: Illustration of our method. The upper part is the overall network architecture, and the lower part is the MPGGD module. The MPGGD generates multiple pseudo-labels $\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$ by multiple different initialization networks ($\{LinkNet(1),...,LinkNet(K)\}$) to fit its M_DSC on true labels $\mathbf{Y}$ to a certain threshold $\varphi$, which subsequently distinguishes the clean pixels set $\mathbf{O}_{c}$ from the noisy pixels set $\mathbf{O}_{n}$ by consistent information.The MPNN is constructed by a model consisting of a student network and a teacher network. Where the student network learns by (i) minimizing the segmentation loss guided by $\mathbf{O}_{c}$ and (ii) minimizing the consistency loss guided by $\mathbf{O}_{n}$ in the teacher model with multiple uncertainties. The teacher network is updated based on the parameters of the student network using the Exponential Moving Average (EMA).
II Methods
To simplify the description of our methodology, we define $\mathbf{D}=\{\mathbf{X}_{(i)},\mathbf{Y}_{(i)}\}_{i=1}^{N}$ to represent the dataset, where $N$ represent the total samples. $\mathbf{X}=\{x_{1},x_{2},...,x_{n}\}$ denotes that the image $\mathbf{X}$ contains $n=w\times h$ pixels and $\mathbf{Y}=\{y_{1},y_{2},...,y_{n}\}$ indicates the corresponding label. Here $y\in\{0,1,2\}$ and the categories in $\{0,1,2\}$ denote the background, optic disc, and optic cup respectively. Fig. 1 illustrates our method. Specifically, the MPGGD generates multiple pseudo-labels $\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$ by multiple different initialization networks ($\{LinkNet(1),...,LinkNet(K)\}$) to fit its M_DSC (Mean DSC of optic disc and cup) on true labels $\mathbf{Y}$ to a certain threshold $\varphi$, which subsequently distinguishes the clean pixels set $\mathbf{O}_{c}$ from the noisy pixels set $\mathbf{O}_{n}$ by consistent information. The MPNN is constructed by a model consisting of a student network and a teacher network. Where the student network learns by (i) minimizing the segmentation loss guided by $\mathbf{O}_{c}$ and (ii) minimizing the consistency loss guided by $\mathbf{O}_{n}$ in the teacher model with multiple uncertainties. The teacher network is updated based on the parameters of the student network using the Exponential Moving Average (EMA).
II-A Multiple Pseudo-labels Generation and Guided Denoising (MPGGD)
Deep learning models can memorize easy samples and then gradually adapt to more difficult instances as the number of training epochs increases. When noisy pixels are present, the deep learning model ends up remembering these incorrect labels, leading to poor generalization performance. Moreover, this phenomenon does not change with training optimization or network architecture choices. Therefore, We contemplate halting the training of the network before it fits the noise and generating pseudo-labels corresponding to the training set. These pseudo-labels are then utilized to distinguish between clean and noisy labels.
In order to distinguish clean pixels from noisy pixels, we use multiple networks with different initializations to fit the same training set, and after M_DSC reaches a certain threshold $\varphi$ stop training to obtain multiple pseudo-labels for the training set. Subsequently, the pixels corresponding to the labels that perform consistently on all the pseudo-labels are considered as clean pixels and the rest are considered as noisy pixels.
Formally, give an image $\mathbf{X}\in\mathbf{D}$, the ground truths are represented $\mathbf{Y}$. We construct a set of $K$ networks $\{LinkNet(1),...,LinkNet(K)\}$ to generate multiple pseudo-labels $\mathbf{Y}_{p}=\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$. All networks are initialized with different values, taking the average dice of the optic disc and cup (M_DSC) as metric. Here $K$ and $\varphi$ are set to 5 and 0.93.
In order to differentiate between noisy pixels and clean pixels, we follow the following process:
$\mathbf{O}_{c}=\{i\mid\mathbf{Y}_{p1}^{i}=\mathbf{Y}_{p2}^{i}=...=\mathbf{Y}_{pK}^{i}\}_{i=0}^{n}.$ (1)
Here, $\mathbf{O}_{c}$ denotes the separated clean pixels set, and the total number of pixels contained in the set of clean pixels is denoted as $s_{c}=|\mathbf{O}_{c}|$. Noisy pixels are those in the entire set excluding clean pixels, and the collection of noisy pixels can be represented as follows:
$\mathbf{O}_{n}=\{i\mid i\notin\ \mathbf{O}_{c}\}_{i=0}^{n}.$ (2)
The total number of pixels contained in the noisy pixels set is denoted as $s_{n}=|\mathbf{O}_{n}|$. These pieces of information will be used to train our network.
TABLE I: Quantitative Results for Different Strategies on the RIGA Test Set. The Test Set Is Set as Average Weight Majority Vote And Individual Rater1
Methods Majority Vote Rater1
$IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
LinkNet [24] 83.86 75.77 91.03 85.64 81.48 75.66 89.66 85.64
Co-teaching [17] 84.76 77.25 91.53 86.58 82.19 76.30 90.08 86.00
TriNet [19] 84.36 77.37 91.29 86.71 82.19 76.83 90.10 86.36
2SRnT [14] 84.50 77.22 91.40 86.62 82.64 76.52 90.34 86.08
MTCL [15] 84.11 76.04 91.15 85.83 82.48 74.85 90.26 85.05
PINT [23] 84.18 77.13 91.22 86.57 82.55 76.11 90.32 85.82
MPNN (Ours) 85.22 78.11 91.83 87.25 83.12 77.34 90.67 86.77
II-B Pixel-level Semi-supervised Pattern Learning
Inherent ambiguity in segmenting objects and background boundaries in the optic disc and cup segmentation task leads to noisy annotations that impact model performance. Therefore, we study the segmentation task of the optic disc and cup with noisy labels and our method draws on the Mean-Teacher [25] model from the classical semi-supervised approach. The student network is parameterized by $\theta$ (updated by back-propagation) and the teacher network is parameterized by $\widetilde{\theta}$ (updated according to the weights $\theta$ at different training stages). The specific computation is the EMA. At training step $t$, $\widetilde{\theta}$ is updated as $\widetilde{\theta}_{t}=\alpha\widetilde{\theta}_{t-1}+(1-\alpha)\theta_{t}$, where $\alpha$ is the EMA decay rate, and it is set to 0.99, as recommended by [25].
For a given image $\mathbf{X}\in\mathbf{D}$, we add different perturbations (Gaussian noise $\{\xi_{1},...,\xi_{m}\}$) to $\mathbf{X}$ to generate a set of inputs $\{\mathbf{X}_{m}\}_{m=1}^{M}$. We input $\mathbf{X}$ into the student network to get the corresponding prediction map $\mathbf{P}_{s}$. At the same time, we input $\{\mathbf{X}_{m}\}_{m=1}^{M}$ into the teacher network to get all prediction probability maps $\{\mathbf{P}_{m}\}_{m=1}^{M}$. Subsequently, the average of all the prediction maps is taken to get the prediction of the teacher network:
$\mathbf{P}_{t}=\frac{1}{M}\sum_{m=1}^{M}{\mathbf{P}_{m}}.$ (3)
The teacher model not only generates goal predictions but also estimates the uncertainty of each goal. Here we follow the approach of [26] and the prediction entropy can be summarized as:
$\begin{split}&u=-\sum_{c}{((\frac{1}{M}\sum_{m=1}^{M}{\mathbf{P}_{m}^{c}}})log{(\frac{1}{M}\sum_{m=1}^{M}{\mathbf{P}_{m}^{c}}))},\end{split}$ (4)
where $\mathbf{P}_{m}^{c}$ is the probability of the $c^{th}$ class in the prediction of $m^{th}$ input. Note that the uncertainty is estimated at the pixel level, and the uncertainty for the entire image $U$ is $\{u\}\in\mathbb{R}^{n}$.
II-B1 Clean Pixels Loss
clean pixels provide the network with reliable and accurate training signals, thus supporting the networks’ ability to learn accurate feature representations and generalization. Our approach extracts a set of clean pixels, enabling the separation of clean and noisy information. Specifically,the clean pixels loss $\mathcal{L}_{c}$ expressed as:
$\mathcal{L}_{c}=\frac{1}{s_{c}}\sum_{i=0}^{s_{c}}{\ell_{ce}(\mathbf{P_{s}}^{\mathbf{O}_{c}^{i}},\mathbf{Y}^{\mathbf{O}_{c}^{i}}}).$ (5)
II-B2 Noisy Pixels Loss
We choose entropy as a metric for estimating uncertainty. When a pixel label tends to be clean, its predictive probability distribution is likely to be peak, indicating low entropy and low uncertainty. Conversely, if a pixel label tends to be noisy, it is likely to have a flatter probability distribution, indicating high entropy and high uncertainty. Therefore, we consider the uncertainty of each pixel as a pixel noise estimate. The set of noisy pixels we separated always contains some easily predictable pixels, and the use of the uncertainty method is to find those clean pixels that can be mistaken as noisy pixels. The consistency loss in the set of noisy pixels is:
$\mathcal{L}_{n}=\frac{\sum_{i=0}^{s_{n}}\mathbbm{I}({u^{\mathbf{O}_{n}^{i}}<H})||\mathbf{P}_{s}^{\mathbf{O}_{n}^{i}}-\mathbf{P}_{t}^{\mathbf{O}_{n}^{i}}||^{2}}{{\textstyle\sum_{i=0}^{s_{n}}}\mathbbm{I}({u^{\mathbf{O}_{n}^{i}}<H})},$ (6)
where $\mathbbm{I}$ is the indicator function and $H$ (initial value is 0.75 and increases with the number of iterations) is a threshold to select the most certain targets. $u^{\mathbf{O}_{n}^{i}}$ is the estimated uncertainty $U$ at $\mathbf{O}_{n}^{i}$ pixel. Note that we only perform the selection of pixels with thresholds higher than $H$ in the noisy pixels set $\mathbf{O}_{n}$.
II-C Final Loss
The total loss function follows the pattern of semi-supervised learning including supervised loss and unsupervised loss. In our approach i.e. clean pixels loss $\mathcal{L}_{c}$ and noisy pixels loss $\mathcal{L}_{n}$. The total loss is expressed as:
$\mathcal{L}_{total}=\alpha\mathcal{L}_{c}+\lambda\mathcal{L}_{n}.$ (7)
Empirically, we set $\alpha=1$ and $\lambda$ is a ramp-up trade-off weight commonly scheduled by the time-dependent Gaussian function [27] $\lambda(t)=w_{max}\cdot e^{(-5(1-\frac{t}{t_{max}})^{2})}$, where $w_{max}$ is the maximum weight commonly set as 0.1 [26] and $t_{max}$ is the maximum training iteration. Such a $\lambda$ weight representation avoids being dominated by misleading targets when starting online training.
III Experiments
III-A Dataset and Pre-processing
The dataset is a publicly available retinal cup and disc segmentation dataset consisting of 750 color fundus images from three different sources: 460 images from MESSIDOR, 195 images from BinRushed, and 95 images from Magrabia. Segmentation masks for the cup and disc contours were manually annotated by six glaucoma specialists following the RIGA benchmarks [28]. During model training, we selected 195 samples from BinRushed and 460 samples from MESSIDOR as the training set as follows [12]. The Magrabia image set containing 95 samples was used as a test set to evaluate the model. We choose the first expert’s annotation as the ground truth for the training set. In order to better evaluate the performance of the method, we construct two test sets, including the mask labeled by the first expert ($\mathbf{Rater1}$) and the mask produced by the majority vote of all experts ($\mathbf{Majority}$ $\mathbf{Vote}$).
III-B Implementation and Evaluation Metrics
We implemented our method and comparative methods in Python using PyTorch and performed the computations on an NVIDIA GeForce RTX 3090 GPU equipped with 8GB of RAM. The backbone network is LinkNet [24] and is trained using the Adam optimizer (betas = (0.9, 0.99)). The initial learning rate was set to 5e-4, divided by 10 for 2000 iterations, with a total of 100 training rounds and an input batch size of 8. The image size of the input network was scaled to 256 × 256 pixels and normalized using the mean and standard deviation of each channel. The model with the best performance on the test set was selected as the final training model. For evaluating the model performance, we give the Dice Similarity Coefficient (DSC) and Intersection over Union (IoU) for each category, excluding the background.
Figure 2: Visualized segmentation results on RIGA of different methods. The ground truth mask and predicted mask are represented by red and green, respectively.
III-C Experiments on RIGA Dataset
Table I presents a comprehensive analysis of the quantitative results obtained from various segmentation methods applied to two distinct test sets (”Majority Vote” and ”Rater1”) within the RIGA dataset. Turning our attention to the performance metrics evaluated, the table showcases results for $IoU_{disc}$, $IoU_{cup}$, $Dice_{disc}$, and $Dice_{cup}$, which serve as key indicators of segmentation accuracy. We observe that LinkNet while delivering competitive performance, falls short of MPNN in terms of all metrics. Co-teaching, TriNet, 2SRnT, MTCL, and PINT also exhibit commendable segmentation results, but they do not surpass the MPNN outcomes. Among the strategies, it’s evident that MPNN emerges as a standout performer. It achieves a remarkable metrics of $IoU_{disc}$ of 85.22%, $IoU_{cup}$ of 78.11%, $Dice_{disc}$ of 91.83%, and $Dice_{cup}$ of 87.25%. Shifting the focus to the ”Rater1” scenario, the trends continue to favor MPNN. Under this scenario, which reflects individual rater annotations, MPNN maintains its performance superiority. With an $IoU_{disc}$ of 83.12%, an $IoU_{cup}$ of 77.34%, a $Dice_{disc}$ of 90.67%, and a $Dice_{cup}$ of 86.77%, MPNN consistently outperforms other methods. In Fig. 2, the visualization results of various methods affirm the efficacy of our proposed approach. These visualizations vividly demonstrate our method’s capacity to precisely segment the targeted regions, aligning seamlessly with the corresponding ground truth annotations. Through these visual representations, we provide qualitative substantiation of our method’s superior performance in contrast to the alternative approaches.
TABLE II: The Ablation Study Of Our Method on RIGA Data. The Impact of Different Numbers of Pseudo-Labels on the Final Result of the Model
Majority Vote
Methods $K$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
3 83.52 76.94 90.81 86.41
4 84.90 77.70 91.62 86.97
5 85.22 78.11 91.83 87.25
MPNN 6 85.19 77.06 91.78 86.53
TABLE III: The Ablation Study Of Our Method on RIGA Data. The Impact of Pseudo-Labels Obtained for Different Threshold Training on the Final Results of the Model
Majority Vote
Methods $\varphi$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
0.91 84.51 76.97 91.39 86.40
0.92 84.73 76.99 91.56 86.47
MPNN 0.93 85.22 78.11 91.83 87.25
0.94 84.79 77.07 91.57 86.51
0.95 83.49 76.68 90.90 86.25
TABLE IV: The Ablation Study Of Our Method on RIGA Data. Comparison of Using Clean Pixels And Noisy Pixels
Methods $\mathbf{O}_{c}$ $\mathbf{O}_{n}$ Majority Vote
$IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
MPNN $\times$ $\times$ 84.28 77.35 91.29 86.70
$\times$ $\checkmark$ 85.12 77.73 91.75 86.94
$\checkmark$ $\times$ 84.97 77.40 91.66 86.74
$\checkmark$ $\checkmark$ 85.22 78.11 91.83 87.25
III-D Parameter Ablation Study
In order to determine suitable parameters for the MPNN method on the RIGA dataset, particularly during the pseudo-label generation process, we conducted ablation experiments. The results of our experiments are shown in Table II and Table III. The evaluation is conducted using the ”Majority Vote” scenario, wherein the obtained results for different metrics.
Table II presents the effect of different numbers of pseudo-labels on the segmentation accuracy of MPNN. When considering the metric $IoU_{disc}$, the model’s performance progressively improves from 83.52% at $K=3$ to 84.90% at $K=4$, and further to a notable 85.22% at $K=5$. Similarly, the metric $IoU_{cup}$ shows improvement from 76.94% at $K=3$ to 77.70% at $K=4$, and then to a substantial 78.11% at $K=5$, signifying an improvement of 1.17% from $K=3$ to $K=5$ in the optic disc and cup segmentation IoU. In addition, the Dice coefficients ($Dice_{disc}$ and $Dice_{cup}$) show similar increases. However, when $K=6$, a noteworthy observation emerges: the performance metrics exhibit a slight decline compared to the results obtained at $K=5$. Specifically, for $IoU_{disc}$, the value drops to 85.19%, a decrease of 0.03% from the value at $K=5$. Similarly, for $IoU_{cup}$, the metric decreases to 77.06%, indicating a reduction of 1.05% compared to the value at $K=5$. This observation underlines the importance of striking a balance when introducing pseudo-labels into the training process. While additional pseudo-labels can aid in capturing more label noise, an excessive quantity might lead to the dilution of valuable prior knowledge. This insight encourages further investigation into the optimal amount of pseudo-labels to integrate, considering the trade-off between information enrichment and the risk of introducing noise.
Figure 3: Visualization of the identified noisy pixels (green), the number underneath indicates the total number of noisy pixels included.
Table III presents the impact of multiple pseudo-labels obtained with different training thresholds $\varphi$ on segmentation accuracy in MPNN. The outcomes indicate that the choice of threshold value during pseudo-label generation has a discernible influence on the model’s performance. Notably, for $IoU_{disc}$, we observe a consistent progression as the threshold value increases. The metric shows an improvement from 84.51% at $\varphi=0.91$ to 84.73% at $\varphi=0.92$ and eventually peaks at an impressive 85.22% for $\varphi=0.93$. However, as $\varphi$ increases to 0.94 and 0.95, $IoU_{disc}$ shows a downward trend. When $\varphi$ increases, the resulting clean pixel set may contain inaccurate information, resulting in reduced model performance. Likewise, $IoU_{cup}$ demonstrates a parallel pattern. The performance rises from 76.97% at $\varphi=0.91$ to 76.99% at $\varphi=0.92$ and subsequently reaches its zenith at 78.11% for $\varphi=0.93$. This signifies a prominent advancement of 1.14% in optic disc and cup segmentation IoU between $\varphi=0.91$ and $\varphi=0.93$. In fact, the selection of the threshold is similar to the selection of the number of pseudo-labels, and the trade-off between rich information and the risk of introducing noise needs to be considered. The results affirm the importance of threshold selection during pseudo-label generation, with the highest threshold value of $\varphi=0.93$ yielding the most favorable outcomes in terms of IoU and Dice scores.
III-E Component Ablation Study
To evaluate the effectiveness of each MPNN component on the RIGA dataset, we performed ablation studies using different variants. The results of our ablation experiments are shown in Table IV. The evaluation is conducted using the ”Majority Vote” scenario, wherein the obtained results for different metrics are presented. Table IV presents the comparison of the performance of the MPNN method between different scenes. The symbol $\times$ indicates that the corresponding pixel set is not distinguished when calculating the corresponding loss.
When we input noise pixels and clean pixels without distinguishing between them, the performance of the MPNN is mediocre. As anticipated, when considering only clean pixels or noisy pixels, the MPNN method has demonstrated significant improvement. Furthermore, when both clean pixels and noisy pixels are utilized simultaneously, the MPNN method further demonstrates enhanced results. Specifically, the $IoU_{disc}$ score stands at 85.22%, showcasing an improvement compared to the use of solely clean pixels. Similarly, $IoU_{cup}$ exhibits an increase to 78.11%. The most significant advancements are observed in $Dice_{disc}$, reaching 91.83%, while $Dice_{cup}$ reaches 87.25%. Fig. 3 shows our isolated collection of noisy pixels, mostly at the segmentation edges. In fact, our strategy for selecting multiple pseudo-labels is a form of the ”overcorrection-is-necessary” strategy. The filtered noisy pixels necessarily contain instances that the model can confidently predict. This also explains why introducing additional noisy labels for consistency loss yields better results.
IV Conclusions
In this paper, we introduce a label-denoising method of the Multiple Pseudo-labels Noise-aware Network (MPNN) for precise optic disc and cup segmentation. To distinguish clean pixels and noisy pixels, The proposed MPGGD module uses multiple networks with different initializations to fit the same training set and stops training after reaching a certain threshold to obtain multiple pseudo-labels of the training set. Subsequently, our method can distinguish between noisy and clean pixels by leveraging consensus among multiple pseudo-labels and it proposes distinct learning strategies for noisy and clean pixels based on teacher-student architecture. Through extensive experimentation on the RIGA dataset, our method performs well on the optic disc and cup segmentation tasks. The visualized outcomes further validate the exceptional performance and robust denoising capabilities of our proposed method. Moreover, our approach exhibits superior segmentation efficacy in optic disc and cup segmentation compared to other label-denoising methods, highlighting our capability to precisely capture object boundaries and generate better masks.

### 15 Mar 2024 v2

@@
-Accurate Segmentation of Optic Disc and Cup from Multiple Pseudo-labels by Noise-Aware Learning
+Accurate Segmentation of Optic Disc and Cup from Multiple Pseudo-labels by Noise-aware Learning thanks: *Corresponding authors (Yang Shen and Shuai Wang)
 Abstract
-Optic disc and cup segmentation play a crucial role in automating the screening and diagnosis of optic glaucoma.
-While data-driven convolutional neural networks (CNNs) show promise in this area, the inherent ambiguity of segmenting object and background boundaries in the task of optic disc and cup segmentation leads to noisy annotations that impact model performance.
+Optic disc and cup segmentation plays a crucial role in automating the screening and diagnosis of optic glaucoma.
+While data-driven convolutional neural networks (CNNs) show promise in this area, the inherent ambiguity of segmenting objects and background boundaries in the task of optic disc and cup segmentation leads to noisy annotations that impact model performance.
 To address this, we propose an innovative label-denoising method of Multiple Pseudo-labels Noise-aware Network (MPNN) for accurate optic disc and cup segmentation.
@@
 The training framework of the MPNN is constructed by a teacher-student architecture to learn segmentation from clean pixels and noisy pixels.
-Particularly, such a framework adeptly leverages (i) reliable and fundamental insights from clean pixels and (ii) the supplementary knowledge within noisy pixels via multiple perturbation-based unsupervised consistency.
-Compared to other label-denoising methods, comprehensive experimental results on the RIGA dataset demonstrate our method’s excellent performance and significant denoising ability.
-The code is available at https://github.com/wwwtttjjj/MPNN
+Particularly, such a framework adeptly leverages (i) reliable and fundamental insight from clean pixels and (ii) the supplementary knowledge within noisy pixels via multiple perturbation-based unsupervised consistency.
+Compared to other label-denoising methods, comprehensive experimental results on the RIGA dataset demonstrate our method’s excellent performance.
+The code is available at https://github.com/wwwtttjjj/MPNN.
 Index Terms:
@@
 With the rapid development of deep learning, various models based on convolutional neural networks (CNNs) have emerged, showcasing promising performance in medical image segmentation.
-These methods [1, 2, 3, 4, 5] leverage their inherent capacity to learn intricate patterns and features from extensive labeled data, enabling them to achieve high levels of accuracy in segmenting medical images.
-However, due to the inherent ambiguity of the true segmentation boundaries in optic disc and cup segmentation [6, 7] and the inevitable errors in manual labeling, the annotation often contains noise.
-Numerous studies [8, 9, 10, 11] have indicated that label noise can significantly impact the accuracy of models.
-In order to mitigate the impact of label noise on medical image segmentation, a strategy is to use multiple expert annotations.
-Some progress has been achieved in the joint annotation learning from multiple experts for the task of optic disc and optic cup segmentation.
+These methods [1, 2, 3] leverage their inherent capacity to learn intricate patterns and features from extensive labeled data, enabling them to achieve high levels of accuracy in segmenting medical images.
+However, due to the inherent ambiguity of the true segmentation boundary in optic disc and cup segmentation and the inevitable errors in manual labeling, the annotation often contains noise.
+Numerous studies [4, 5] have indicated that label noise can significantly impact the accuracy of models.
+To mitigate the impact of label noise on medical image segmentation, one strategy is to use multiple expert annotations.
 Ji et al.
-[12] introduced MRNet, an innovative technique that capitalizes on the specialized insights of individual raters, treating their expertise as prior knowledge to extract high-level semantic features.
-Their proposed approach involves reconstructing multi-rater rankings based on initial predictions, subsequently harnessing the varying degrees of agreement among multiple raters to enhance the overall segmentation performance.
-By amalgamating expert knowledge and effectively pooling input from multiple raters, MRNet presents a promising avenue for improving the precision of optic disc and cup segmentation.
-Self-Calib [13] attempts to find ground truths by running divergent and convergent models in loops.
-The latter category of methods demonstrates a greater capacity for exploiting information from multiple annotations, resulting in a more impartial segmentation model.
-However, it is worth noting that obtaining segmentation annotations from experts is a challenging task, not to mention the need for annotations from multiple experts.
+[6] introduced MRNet, an innovative technique that capitalizes on the specialized insight of individual raters, treating their expertise as prior knowledge to extract high-level semantic features.
+Self-Calib [7] attempts to find ground truths by running divergent and convergent models in loops, demonstrating a greater capacity for exploiting information from multiple annotations, which results in a more impartial segmentation model.
+However, it is worth noting that obtaining segmentation annotations from multiple experts is a challenging task.
 Therefore, it is necessary to explore other available methods.
 Without introducing additional expert labels, a series of label-denoising strategies have been proposed to cope with the challenges posed by noisy labels in segmentation tasks.
-These strategies typically involve two steps: first, noisy labels are identified using various methods, and then pseudo-labels are used to refine these noisy labels or to reject them.
-Some methods [8, 14, 15] use Confident Learning [16] modules to find noisy labels and use refinement module for processing.
-The method based on Confidence Learning assesses the classification confidence of each pixel’s output from the network and identifies potential noisy pixels based on this confidence level.
-Subsequently, the refinement module uses pseudo-labels to replace these noisy labels.
-Other methods [17, 18, 19, 20, 21, 22, 23] involve training multiple networks and selecting labels with the minimum loss consensus between network predictions for training.
-In essence, these selected pixels are treated as clean pixels, thereby eliminating noisy pixels.
-Compared to multi-annotation learning, the direct denoising approach is simple, efficient, and does not require the introduction of additional expert knowledge.
-However, while these methods have shown promise in various domains, their adaptation and implementation for the specific task of optic disc and cup segmentation still require careful investigation.
+These strategies typically involve two steps: first, noisy labels are identified using various methods, and then there is the option of discarding them or refining them using pseudo-labels.
+Some methods [8, 9, 10, 11] involve training multiple networks and selecting labels with the minimum loss consensus between network predictions for training.
+Other methods [4, 12, 13] use Confident Learning [14] modules to find noisy labels and use refinement module for processing.
+However, their research often involves artificially simulated noisy labels and annotations from non-experts.
+There has not been specific research on annotation noise caused by ambiguous boundaries in the segmentation of optic disc and cup.
 In this paper, we propose an innovative label-denoising method of Multiple Pseudo-labels Noise-aware Network (MPNN) for accurate optic disc and cup segmentation.
@@
 We use multiple networks with different initialization to fit the same training set, and after reaching a certain threshold stop training to obtain multiple pseudo-labels for the training set.
-Subsequently, the pixels corresponding to the labels that perform consistently on all the pseudo-labels are considered as clean pixels and the rest are considered as noisy pixels.
+Subsequently, the pixels corresponding to the labels that perform consistently on all the pseudo-labels are considered clean pixels, and the rest are considered noisy pixels.
 Furthermore, the proposed MPNN consists of a teacher-student architecture, where the student network learns by (i) minimizing the segmentation loss on the set of clean pixels and (ii) minimizing the consistency loss on the set of noisy pixels in the teacher model with multiple uncertainties.
@@
   • $\bullet$
-    To the best of our knowledge, the proposed MPNN is the first method to improve the accuracy of the optic disc and cup from the perspective of label denoising.
+    To the best of our knowledge, the proposed MPNN is the first method to improve the segmentation accuracy of the optic disc and cup from the perspective of label denoising.
   • $\bullet$
-    We proposed the MPGGD as the first execution module.
-It separates reliable and unreliable information, provides accurate priors for the segmentation network, and avoids the negative impact of label noise on the network.
+    We propose the MPGGD module, which separates reliable and unreliable information, provides accurate prior information for the network and avoids the negative impact of label noise on the network.
 Figure 1: Illustration of our method.
-The upper part is the overall network architecture, and the lower part is the MPGGD module.
-The MPGGD generates multiple pseudo-labels $\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$ by multiple different initialization networks ($\{LinkNet(1),...,LinkNet(K)\}$) to fit its M_DSC on true labels $\mathbf{Y}$ to a certain threshold $\varphi$, which subsequently distinguishes the clean pixels set $\mathbf{O}_{c}$ from the noisy pixels set $\mathbf{O}_{n}$ by consistent information.The MPNN is constructed by a model consisting of a student network and a teacher network.
-Where the student network learns by (i) minimizing the segmentation loss guided by $\mathbf{O}_{c}$ and (ii) minimizing the consistency loss guided by $\mathbf{O}_{n}$ in the teacher model with multiple uncertainties.
-The teacher network is updated based on the parameters of the student network using the Exponential Moving Average (EMA).
+(a) the architecture of MPNN; (b) the MPGGD module.
+The MPGGD generates multiple pseudo-labels $\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$ by multiple different initialization networks ($\{LinkNet(1),...,LinkNet(K)\}$) to fit its $DSC_{m}$ on true label $\mathbf{Y}$ to a certain threshold $\varphi$, which subsequently distinguishes the clean pixels set $\mathbf{O}_{cl}$ from the noisy pixels set $\mathbf{O}_{no}$ by consistent information.
+The MPNN is constructed by a model consisting of a student network and a teacher network.
+The student network learns by (i) minimizing the segmentation loss guided by $\mathbf{O}_{cl}$ and (ii) minimizing the consistency loss guided by $\mathbf{O}_{no}$ in the teacher model with multiple uncertainties.
+The teacher network is updated using the EMA.
 II Methods
 To simplify the description of our methodology, we define $\mathbf{D}=\{\mathbf{X}_{(i)},\mathbf{Y}_{(i)}\}_{i=1}^{N}$ to represent the dataset, where $N$ represent the total samples.
-$\mathbf{X}=\{x_{1},x_{2},...,x_{n}\}$ denotes that the image $\mathbf{X}$ contains $n=w\times h$ pixels and $\mathbf{Y}=\{y_{1},y_{2},...,y_{n}\}$ indicates the corresponding label.
+$\mathbf{X}=\{x_{1},x_{2},...,x_{n}\}$ denotes that the image $\mathbf{X}$ contains $n=w\times h$ pixels and $\mathbf{Y}=\{y_{1},y_{2},...,y_{n}\}$ indicates the corresponding labels.
 Here $y\in\{0,1,2\}$ and the categories in $\{0,1,2\}$ denote the background, optic disc, and optic cup respectively.
 Fig.
-1 illustrates our method.
-Specifically, the MPGGD generates multiple pseudo-labels $\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$ by multiple different initialization networks ($\{LinkNet(1),...,LinkNet(K)\}$) to fit its M_DSC (Mean DSC of optic disc and cup) on true labels $\mathbf{Y}$ to a certain threshold $\varphi$, which subsequently distinguishes the clean pixels set $\mathbf{O}_{c}$ from the noisy pixels set $\mathbf{O}_{n}$ by consistent information.
-The MPNN is constructed by a model consisting of a student network and a teacher network.
-Where the student network learns by (i) minimizing the segmentation loss guided by $\mathbf{O}_{c}$ and (ii) minimizing the consistency loss guided by $\mathbf{O}_{n}$ in the teacher model with multiple uncertainties.
-The teacher network is updated based on the parameters of the student network using the Exponential Moving Average (EMA).
+1 illustrates the method of MPNN.
+More details will be described in subsequent sections.
 II-A Multiple Pseudo-labels Generation and Guided Denoising (MPGGD)
 Deep learning models can memorize easy samples and then gradually adapt to more difficult instances as the number of training epochs increases.
-When noisy pixels are present, the deep learning model ends up remembering these incorrect labels, leading to poor generalization performance.
-Moreover, this phenomenon does not change with training optimization or network architecture choices.
-Therefore, We contemplate halting the training of the network before it fits the noise and generating pseudo-labels corresponding to the training set.
-These pseudo-labels are then utilized to distinguish between clean and noisy labels.
-In order to distinguish clean pixels from noisy pixels, we use multiple networks with different initializations to fit the same training set, and after M_DSC reaches a certain threshold $\varphi$ stop training to obtain multiple pseudo-labels for the training set.
-Subsequently, the pixels corresponding to the labels that perform consistently on all the pseudo-labels are considered as clean pixels and the rest are considered as noisy pixels.
-Formally, give an image $\mathbf{X}\in\mathbf{D}$, the ground truths are represented $\mathbf{Y}$.
+When noisy labels are present, the deep learning model ends up remembering these incorrect labels, leading to poor generalization performance.
+Therefore, we contemplate halting the training of the network before it fits the noise and generating multiple pseudo-labels corresponding to the training set.
+These pseudo-labels are utilized to distinguish between clean and noisy pixels.
+Specifically, to distinguish clean pixels from noisy pixels, we use multiple networks with different initializations to fit the same training set, and after $DSC_{m}$ (Mean Dice Similarity Coefficient of the optic disc and cup) reaches a certain threshold $\varphi$ stop training to obtain multiple pseudo-labels of the training set.
+To maximize the identification of annotation noise, we adopt an over-correction strategy, that is, pixels corresponding to labels that behave consistently across all pseudo-labels are considered clean pixels, and the rest are considered noisy pixels.
+Formally, give an image $\mathbf{X}\in\mathbf{D}$, the ground truth is represented $\mathbf{Y}$.
 We construct a set of $K$ networks $\{LinkNet(1),...,LinkNet(K)\}$ to generate multiple pseudo-labels $\mathbf{Y}_{p}=\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$.
-All networks are initialized with different values, taking the average dice of the optic disc and cup (M_DSC) as metric.
-Here $K$ and $\varphi$ are set to 5 and 0.93.
-In order to differentiate between noisy pixels and clean pixels, we follow the following process:
-$\mathbf{O}_{c}=\{i\mid\mathbf{Y}_{p1}^{i}=\mathbf{Y}_{p2}^{i}=...=\mathbf{Y}_{pK}^{i}\}_{i=0}^{n}.$ (1)
-Here, $\mathbf{O}_{c}$ denotes the separated clean pixels set, and the total number of pixels contained in the set of clean pixels is denoted as $s_{c}=|\mathbf{O}_{c}|$.
-Noisy pixels are those in the entire set excluding clean pixels, and the collection of noisy pixels can be represented as follows:
-$\mathbf{O}_{n}=\{i\mid i\notin\ \mathbf{O}_{c}\}_{i=0}^{n}.$ (2)
-The total number of pixels contained in the noisy pixels set is denoted as $s_{n}=|\mathbf{O}_{n}|$. These pieces of information will be used to train our network.
-TABLE I: Quantitative Results for Different Strategies on the RIGA Test Set. The Test Set Is Set as Average Weight Majority Vote And Individual Rater1
-Methods Majority Vote Rater1
-$IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
-LinkNet [24] 83.86 75.77 91.03 85.64 81.48 75.66 89.66 85.64
-Co-teaching [17] 84.76 77.25 91.53 86.58 82.19 76.30 90.08 86.00
-TriNet [19] 84.36 77.37 91.29 86.71 82.19 76.83 90.10 86.36
-2SRnT [14] 84.50 77.22 91.40 86.62 82.64 76.52 90.34 86.08
-MTCL [15] 84.11 76.04 91.15 85.83 82.48 74.85 90.26 85.05
-PINT [23] 84.18 77.13 91.22 86.57 82.55 76.11 90.32 85.82
-MPNN (Ours) 85.22 78.11 91.83 87.25 83.12 77.34 90.67 86.77
+All networks are initialized with different values, taking the $DSC_{m}$ as the metric.
+Considering the scarcity of annotation noise, here $K$ and $\varphi$ are set to 5 and 0.93.
+To differentiate between clean pixels and noisy pixels, we follow the following process:
+$\mathbf{O}_{cl}=\{i\mid\mathbf{Y}_{p1}^{i}=\mathbf{Y}_{p2}^{i}=...=\mathbf{Y}_{pK}^{i}\}_{i=0}^{n}.$ (1)
+Here, $\mathbf{O}_{cl}$ denotes the separated clean pixels set, and the total number of pixels contained in the set of clean pixels is denoted as $s_{cl}=|\mathbf{O}_{cl}|$.
+Obviously, the noisy pixels set can be represented as follows:
+$\mathbf{O}_{no}=\{i\mid i\notin\ \mathbf{O}_{cl}\}_{i=0}^{n}.$ (2)
+The total number of pixels contained in the noisy pixels set is denoted as $s_{no}=|\mathbf{O}_{no}|$. These pieces of information will be used to guide the training of the MPNN.
 II-B Pixel-level Semi-supervised Pattern Learning
 Inherent ambiguity in segmenting objects and background boundaries in the optic disc and cup segmentation task leads to noisy annotations that impact model performance.
-Therefore, we study the segmentation task of the optic disc and cup with noisy labels and our method draws on the Mean-Teacher [25] model from the classical semi-supervised approach.
+Therefore, we study the segmentation task of optic disc and cup with noisy labels and our method draws on the Mean-Teacher [15] model from the classical semi-supervised approach.
 The student network is parameterized by $\theta$ (updated by back-propagation) and the teacher network is parameterized by $\widetilde{\theta}$ (updated according to the weights $\theta$ at different training stages).
@@
 Our approach extracts a set of clean pixels, enabling the separation of clean and noisy information.
-Specifically,the clean pixels loss $\mathcal{L}_{c}$ expressed as:
-$\mathcal{L}_{c}=\frac{1}{s_{c}}\sum_{i=0}^{s_{c}}{\ell_{ce}(\mathbf{P_{s}}^{\mathbf{O}_{c}^{i}},\mathbf{Y}^{\mathbf{O}_{c}^{i}}}).$ (5)
+Specifically, the clean pixels loss $\mathcal{L}_{cl}$ expressed as:
+$\mathcal{L}_{cl}=\frac{1}{s_{cl}}\sum_{i=0}^{s_{cl}}{\ell_{ce}(\mathbf{P_{s}}^{\mathbf{O}_{cl}^{i}},\mathbf{Y}^{\mathbf{O}_{cl}^{i}}}).$ (5)
+TABLE I: Comparative Results of Different Strategies on the Majority Vote Test Set and the Rater1 Test Set, Respectively
+Methods Majority Vote Rater1
+$IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
+LinkNet [17] 83.86 75.77 91.03 85.64 81.48 75.66 89.66 85.64
+Co-teaching [8] 84.76 77.25 91.53 86.58 82.19 76.30 90.08 86.00
+TriNet [10] 84.36 77.37 91.29 86.71 82.19 76.83 90.10 86.36
+2SRnT [12] 84.50 77.22 91.40 86.62 82.64 76.52 90.34 86.08
+MTCL [13] 84.11 76.04 91.15 85.83 82.48 74.85 90.26 85.05
+PINT [11] 84.18 77.13 91.22 86.57 82.55 76.11 90.32 85.82
+MPNN (Ours) 85.22 78.11 91.83 87.25 83.12 77.34 90.67 86.77
 II-B2 Noisy Pixels Loss
@@
 Therefore, we consider the uncertainty of each pixel as a pixel noise estimate.
-The set of noisy pixels we separated always contains some easily predictable pixels, and the use of the uncertainty method is to find those clean pixels that can be mistaken as noisy pixels.
-The consistency loss in the set of noisy pixels is:
-$\mathcal{L}_{n}=\frac{\sum_{i=0}^{s_{n}}\mathbbm{I}({u^{\mathbf{O}_{n}^{i}}<H})||\mathbf{P}_{s}^{\mathbf{O}_{n}^{i}}-\mathbf{P}_{t}^{\mathbf{O}_{n}^{i}}||^{2}}{{\textstyle\sum_{i=0}^{s_{n}}}\mathbbm{I}({u^{\mathbf{O}_{n}^{i}}<H})},$ (6)
+Considering the MPGGD method is an over-correction strategy, the set of noisy pixels we separated always contains some easily predictable pixels, and the use of the uncertainty method is to find those clean pixels that can be mistaken as noisy pixels.
+The consistency loss $\mathcal{L}_{no}$ in the set of noisy pixels expressed as:
+$\mathcal{L}_{no}=\frac{\sum_{i=0}^{s_{no}}\mathbbm{I}({u^{\mathbf{O}_{no}^{i}}<H})||\mathbf{P}_{s}^{\mathbf{O}_{no}^{i}}-\mathbf{P}_{t}^{\mathbf{O}_{no}^{i}}||^{2}}{{\textstyle\sum_{i=0}^{s_{no}}}\mathbbm{I}({u^{\mathbf{O}_{no}^{i}}<H})},$ (6)
 where $\mathbbm{I}$ is the indicator function and $H$ (initial value is 0.75 and increases with the number of iterations) is a threshold to select the most certain targets.
-$u^{\mathbf{O}_{n}^{i}}$ is the estimated uncertainty $U$ at $\mathbf{O}_{n}^{i}$ pixel.
-Note that we only perform the selection of pixels with thresholds higher than $H$ in the noisy pixels set $\mathbf{O}_{n}$.
+$u^{\mathbf{O}_{no}^{i}}$ is the estimated uncertainty $U$ at $\mathbf{O}_{no}^{i}$ pixel.
+Note that we only perform the selection of pixels with thresholds higher than $H$ in the noisy pixels set $\mathbf{O}_{no}$.
+Figure 2: Visualized segmentation results on the RIGA dataset of different methods. The ground truth mask and predicted mask are represented by red and green, respectively.
 II-C Final Loss
-The total loss function follows the pattern of semi-supervised learning including supervised loss and unsupervised loss.
-In our approach i.e.
-clean pixels loss $\mathcal{L}_{c}$ and noisy pixels loss $\mathcal{L}_{n}$.
-The total loss is expressed as:
-$\mathcal{L}_{total}=\alpha\mathcal{L}_{c}+\lambda\mathcal{L}_{n}.$ (7)
-Empirically, we set $\alpha=1$ and $\lambda$ is a ramp-up trade-off weight commonly scheduled by the time-dependent Gaussian function [27] $\lambda(t)=w_{max}\cdot e^{(-5(1-\frac{t}{t_{max}})^{2})}$, where $w_{max}$ is the maximum weight commonly set as 0.1 [26] and $t_{max}$ is the maximum training iteration.
+The total loss function follows the pattern of semi-supervised learning including supervised loss $\mathcal{L}_{cl}$ and unsupervised loss $\mathcal{L}_{no}$. The total loss is expressed as:
+$\mathcal{L}_{total}=\beta\mathcal{L}_{cl}+\lambda\mathcal{L}_{no}.$ (7)
+Empirically, we set $\beta=1$ and $\lambda$ is a ramp-up trade-off weight commonly scheduled by the time-dependent Gaussian function [18] $\lambda(t)=w_{max}\cdot e^{(-5(1-\frac{t}{t_{max}})^{2})}$, where $w_{max}$ is the maximum weight commonly set as 0.1 [16] and $t_{max}$ is the maximum training iteration.
 Such a $\lambda$ weight representation avoids being dominated by misleading targets when starting online training.
@@
 III-A Dataset and Pre-processing
-The dataset is a publicly available retinal cup and disc segmentation dataset consisting of 750 color fundus images from three different sources: 460 images from MESSIDOR, 195 images from BinRushed, and 95 images from Magrabia.
-Segmentation masks for the cup and disc contours were manually annotated by six glaucoma specialists following the RIGA benchmarks [28].
-During model training, we selected 195 samples from BinRushed and 460 samples from MESSIDOR as the training set as follows [12].
+The RIGA benchmark [19] is a publicly available retinal disc and cup segmentation dataset consisting of 750 color fundus images from three different sources: 460 images from MESSIDOR, 195 images from BinRushed, and 95 images from Magrabia.
+Segmentation masks for the disc and cup contours were manually annotated by six glaucoma specialists following the RIGA benchmark [19].
+We selected 195 samples from BinRushed and 460 samples from MESSIDOR as the training set as follows [6].
 The Magrabia image set containing 95 samples was used as a test set to evaluate the model.
 We choose the first expert’s annotation as the ground truth for the training set.
-In order to better evaluate the performance of the method, we construct two test sets, including the mask labeled by the first expert ($\mathbf{Rater1}$) and the mask produced by the majority vote of all experts ($\mathbf{Majority}$ $\mathbf{Vote}$).
+To better evaluate the performance of the method, we construct two test sets, using the masks labeled by the first expert ($\mathbf{Rater1}$) and produced by the majority vote of all experts ($\mathbf{Majority}$ $\mathbf{Vote}$) as ground truth, respectively.
 III-B Implementation and Evaluation Metrics
-We implemented our method and comparative methods in Python using PyTorch and performed the computations on an NVIDIA GeForce RTX 3090 GPU equipped with 8GB of RAM.
-The backbone network is LinkNet [24] and is trained using the Adam optimizer (betas = (0.9, 0.99)).
-The initial learning rate was set to 5e-4, divided by 10 for 2000 iterations, with a total of 100 training rounds and an input batch size of 8.
-The image size of the input network was scaled to 256 × 256 pixels and normalized using the mean and standard deviation of each channel.
-The model with the best performance on the test set was selected as the final training model.
-For evaluating the model performance, we give the Dice Similarity Coefficient (DSC) and Intersection over Union (IoU) for each category, excluding the background.
-Figure 2: Visualized segmentation results on RIGA of different methods. The ground truth mask and predicted mask are represented by red and green, respectively.
-III-C Experiments on RIGA Dataset
-Table I presents a comprehensive analysis of the quantitative results obtained from various segmentation methods applied to two distinct test sets (”Majority Vote” and ”Rater1”) within the RIGA dataset.
+We implemented our method and comparative methods in Python using PyTorch and performed the computations on an NVIDIA GeForce RTX 3090 GPU equipped with 24GB of RAM.
+We utilized the LinkNet [17] as the backbone and trained it using the Adam optimizer (betas = (0.9, 0.99)).
+We set the initial learning rate to 5e-4, decreased it by a factor of 10 every 2000 iterations, conducted a total of 100 training rounds, and maintained a batch size of 8.
+The input network’s image size was scaled to 256 × 256 pixels and normalized using the mean and standard deviation for each channel.
+To evaluate the model’s performance, we calculated the Intersection over Union (IoU) and Dice for each category, excluding the background.
+III-C Experiments on the RIGA Dataset
+Table I presents the quantitative results obtained from various segmentation methods applied to two distinct test sets ($\mathbf{Majority}$ $\mathbf{Vote}$ and $\mathbf{Rater1}$).
 Turning our attention to the performance metrics evaluated, the table showcases results for $IoU_{disc}$, $IoU_{cup}$, $Dice_{disc}$, and $Dice_{cup}$, which serve as key indicators of segmentation accuracy.
-We observe that LinkNet while delivering competitive performance, falls short of MPNN in terms of all metrics.
-Co-teaching, TriNet, 2SRnT, MTCL, and PINT also exhibit commendable segmentation results, but they do not surpass the MPNN outcomes.
+We observe that LinkNet while delivering competitive performance, falls short of the MPNN in terms of all metrics.
+Co-teaching, TriNet, 2SRnT, MTCL, and PINT also exhibit commendable results, but they do not surpass the MPNN outcomes.
 Among the strategies, it’s evident that MPNN emerges as a standout performer.
 It achieves a remarkable metrics of $IoU_{disc}$ of 85.22%, $IoU_{cup}$ of 78.11%, $Dice_{disc}$ of 91.83%, and $Dice_{cup}$ of 87.25%.
-Shifting the focus to the ”Rater1” scenario, the trends continue to favor MPNN.
-Under this scenario, which reflects individual rater annotations, MPNN maintains its performance superiority.
-With an $IoU_{disc}$ of 83.12%, an $IoU_{cup}$ of 77.34%, a $Dice_{disc}$ of 90.67%, and a $Dice_{cup}$ of 86.77%, MPNN consistently outperforms other methods.
+Shifting the focus to the scenario of $\mathbf{Rater1}$, the trends continue to favor MPNN.
 In Fig.
-2, the visualization results of various methods affirm the efficacy of our proposed approach.
-These visualizations vividly demonstrate our method’s capacity to precisely segment the targeted regions, aligning seamlessly with the corresponding ground truth annotations.
-Through these visual representations, we provide qualitative substantiation of our method’s superior performance in contrast to the alternative approaches.
-TABLE II: The Ablation Study Of Our Method on RIGA Data. The Impact of Different Numbers of Pseudo-Labels on the Final Result of the Model
+2, the visualization results of various methods affirm the efficacy of the MPNN.
+These visualizations demonstrate our method’s capacity to precisely segment the targeted regions.
+TABLE II: Ablation Study of Our Method on the Majority Vote Test Set. The Impact of Different Numbers of Pseudo-Labels on the Final Results
 Majority Vote
-Methods $K$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
+Method $K$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
 3 83.52 76.94 90.81 86.41
@@
 MPNN 6 85.19 77.06 91.78 86.53
-TABLE III: The Ablation Study Of Our Method on RIGA Data. The Impact of Pseudo-Labels Obtained for Different Threshold Training on the Final Results of the Model
+TABLE III: Ablation Study of Our Method on the Majority Vote Test Set. The Impact of Pseudo-Labels Obtained for Different Threshold Training on the Final Results
 Majority Vote
-Methods $\varphi$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
+Method $\varphi$ $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
 0.91 84.51 76.97 91.39 86.40
@@
 0.95 83.49 76.68 90.90 86.25
-TABLE IV: The Ablation Study Of Our Method on RIGA Data. Comparison of Using Clean Pixels And Noisy Pixels
-Methods $\mathbf{O}_{c}$ $\mathbf{O}_{n}$ Majority Vote
+III-D Parameter Ablation Study
+To determine suitable parameters for the MPNN on the RIGA dataset, particularly during the pseudo-labels generation process, we conducted ablation experiments in the scenario of $\mathbf{Majority}$ $\mathbf{Vote}$.
+Table II presents the effect of different numbers of pseudo-labels on the segmentation accuracy of the MPNN.
+When considering the metric $IoU_{disc}$, the model’s performance progressively improves from 83.52% at $K=3$ to 84.90% at $K=4$, and further to a notable 85.22% at $K=5$.
+Similarly, the metric $IoU_{cup}$ shows improvement from 76.94% at $K=3$ to 77.70% at $K=4$, and then to a substantial 78.11% at $K=5$.
+In addition, the Dice coefficients ($Dice_{disc}$ and $Dice_{cup}$) show similar increases.
+However, when $K=6$, a noteworthy observation emerges: the performance metrics exhibit a slight decline compared to the results obtained at $K=5$.
+This observation underlines the importance of striking a balance when introducing pseudo-labels into the training process.
+While additional pseudo-labels can aid in capturing more label noise, an excessive quantity might lead to the dilution of valuable prior knowledge.
+This insight encourages further investigation into the optimal amount of pseudo-labels to integrate, considering the trade-off between information enrichment and the risk of introducing noise.
+Table III presents the impact of multiple pseudo-labels obtained with different training thresholds $\varphi$ on the segmentation accuracy of MPNN.
+Notably, for $IoU_{disc}$, we observe a consistent progression as the threshold value increases.
+The metric shows an improvement from 84.51% at $\varphi=0.91$ to 84.73% at $\varphi=0.92$ and eventually peaks at an impressive 85.22% for $\varphi=0.93$.
+However, as $\varphi$ increases to 0.94 and 0.95, $IoU_{disc}$ shows a downward trend.
+When $\varphi$ increases, the resulting clean pixel set may contain inaccurate information, resulting in reduced model performance.
+The results affirm the importance of threshold selection during pseudo-label generation, with the highest threshold value of $\varphi=0.93$ yielding the most favorable outcomes in terms of IoU and Dice scores.
+TABLE IV: Ablation Study of Our Method on the Majority Vote Test Set. Comparison Results of Using Clean Pixels And Noisy Pixels
+Method $\mathbf{O}_{cl}$ $\mathbf{O}_{no}$ Majority Vote
 $IoU_{disc}(\%)\uparrow$ $IoU_{cup}(\%)\uparrow$ $Dice_{disc}(\%)\uparrow$ $Dice_{cup}(\%)\uparrow$
@@
 $\checkmark$ $\checkmark$ 85.22 78.11 91.83 87.25
-III-D Parameter Ablation Study
-In order to determine suitable parameters for the MPNN method on the RIGA dataset, particularly during the pseudo-label generation process, we conducted ablation experiments.
-The results of our experiments are shown in Table II and Table III.
-The evaluation is conducted using the ”Majority Vote” scenario, wherein the obtained results for different metrics.
-Table II presents the effect of different numbers of pseudo-labels on the segmentation accuracy of MPNN.
-When considering the metric $IoU_{disc}$, the model’s performance progressively improves from 83.52% at $K=3$ to 84.90% at $K=4$, and further to a notable 85.22% at $K=5$.
-Similarly, the metric $IoU_{cup}$ shows improvement from 76.94% at $K=3$ to 77.70% at $K=4$, and then to a substantial 78.11% at $K=5$, signifying an improvement of 1.17% from $K=3$ to $K=5$ in the optic disc and cup segmentation IoU.
-In addition, the Dice coefficients ($Dice_{disc}$ and $Dice_{cup}$) show similar increases.
-However, when $K=6$, a noteworthy observation emerges: the performance metrics exhibit a slight decline compared to the results obtained at $K=5$.
-Specifically, for $IoU_{disc}$, the value drops to 85.19%, a decrease of 0.03% from the value at $K=5$.
-Similarly, for $IoU_{cup}$, the metric decreases to 77.06%, indicating a reduction of 1.05% compared to the value at $K=5$.
-This observation underlines the importance of striking a balance when introducing pseudo-labels into the training process.
-While additional pseudo-labels can aid in capturing more label noise, an excessive quantity might lead to the dilution of valuable prior knowledge.
-This insight encourages further investigation into the optimal amount of pseudo-labels to integrate, considering the trade-off between information enrichment and the risk of introducing noise.
-Figure 3: Visualization of the identified noisy pixels (green), the number underneath indicates the total number of noisy pixels included.
-Table III presents the impact of multiple pseudo-labels obtained with different training thresholds $\varphi$ on segmentation accuracy in MPNN.
-The outcomes indicate that the choice of threshold value during pseudo-label generation has a discernible influence on the model’s performance.
-Notably, for $IoU_{disc}$, we observe a consistent progression as the threshold value increases.
-The metric shows an improvement from 84.51% at $\varphi=0.91$ to 84.73% at $\varphi=0.92$ and eventually peaks at an impressive 85.22% for $\varphi=0.93$.
-However, as $\varphi$ increases to 0.94 and 0.95, $IoU_{disc}$ shows a downward trend.
-When $\varphi$ increases, the resulting clean pixel set may contain inaccurate information, resulting in reduced model performance.
-Likewise, $IoU_{cup}$ demonstrates a parallel pattern.
-The performance rises from 76.97% at $\varphi=0.91$ to 76.99% at $\varphi=0.92$ and subsequently reaches its zenith at 78.11% for $\varphi=0.93$.
-This signifies a prominent advancement of 1.14% in optic disc and cup segmentation IoU between $\varphi=0.91$ and $\varphi=0.93$.
-In fact, the selection of the threshold is similar to the selection of the number of pseudo-labels, and the trade-off between rich information and the risk of introducing noise needs to be considered.
-The results affirm the importance of threshold selection during pseudo-label generation, with the highest threshold value of $\varphi=0.93$ yielding the most favorable outcomes in terms of IoU and Dice scores.
 III-E Component Ablation Study
-To evaluate the effectiveness of each MPNN component on the RIGA dataset, we performed ablation studies using different variants.
+To evaluate the effectiveness of each MPNN component, we performed ablation studies using different variants.
 The results of our ablation experiments are shown in Table IV.
-The evaluation is conducted using the ”Majority Vote” scenario, wherein the obtained results for different metrics are presented.
-Table IV presents the comparison of the performance of the MPNN method between different scenes.
+The evaluation is conducted in the scenario of $\mathbf{Majority}$ $\mathbf{Vote}$.
 The symbol $\times$ indicates that the corresponding pixel set is not distinguished when calculating the corresponding loss.
-When we input noise pixels and clean pixels without distinguishing between them, the performance of the MPNN is mediocre.
+When we input noisy pixels and clean pixels without distinguishing between them, the performance of the MPNN is mediocre.
 As anticipated, when considering only clean pixels or noisy pixels, the MPNN method has demonstrated significant improvement.
 Furthermore, when both clean pixels and noisy pixels are utilized simultaneously, the MPNN method further demonstrates enhanced results.
-Specifically, the $IoU_{disc}$ score stands at 85.22%, showcasing an improvement compared to the use of solely clean pixels.
-Similarly, $IoU_{cup}$ exhibits an increase to 78.11%.
-The most significant advancements are observed in $Dice_{disc}$, reaching 91.83%, while $Dice_{cup}$ reaches 87.25%.
 Fig.
 3 shows our isolated collection of noisy pixels, mostly at the segmentation edges.
-In fact, our strategy for selecting multiple pseudo-labels is a form of the ”overcorrection-is-necessary” strategy.
+Our strategy for selecting multiple pseudo-labels is a form of the “overcorrection-is-necessary” strategy.
 The filtered noisy pixels necessarily contain instances that the model can confidently predict.
 This also explains why introducing additional noisy labels for consistency loss yields better results.
+Figure 3: Visualization of the identified noisy pixels (green), the number underneath indicates the total number of noisy pixels included.
 IV Conclusions
 In this paper, we introduce a label-denoising method of the Multiple Pseudo-labels Noise-aware Network (MPNN) for precise optic disc and cup segmentation.
-To distinguish clean pixels and noisy pixels, The proposed MPGGD module uses multiple networks with different initializations to fit the same training set and stops training after reaching a certain threshold to obtain multiple pseudo-labels of the training set.
+The proposed MPGGD module uses multiple networks with different initializations to fit the same training set and stops training after reaching a certain threshold to obtain multiple pseudo-labels of the training set.
 Subsequently, our method can distinguish between noisy and clean pixels by leveraging consensus among multiple pseudo-labels and it proposes distinct learning strategies for noisy and clean pixels based on teacher-student architecture.
@@
 The visualized outcomes further validate the exceptional performance and robust denoising capabilities of our proposed method.
-Moreover, our approach exhibits superior segmentation efficacy in optic disc and cup segmentation compared to other label-denoising methods, highlighting our capability to precisely capture object boundaries and generate better masks.
+However, there are significant differences among various medical image segmentation tasks with ambiguous boundaries, and such variations can noticeably impact algorithm performance.
+Therefore, the successful extension of the proposed MPNN to other domains of medical image segmentation with ambiguous boundaries is a direction worthy of in-depth exploration.
+V Acknowledge
+This research was supported by the Zhejiang Provincial Natural Science Foundation of China under Grant No.
+LDT23F01015F01 and No.
+LY21F020004, Natural Science Foundation of Jiangsu Province under Grant No.
+BK20220266, and National Natural Science Foundation of China No.
+61572243 and Construction Project of Lishui University Discipline (Zhejiang Province First Class Discipline, Discipline name: intelligent science and technology) No.
+XK0430403005.